A bag contains 5 green, 3 red, 2 orange and 6 yellow skittles. How many green skittles must be added to the bag so that the probability of ramdonly picking a green skittle is 2/3?
Answer by stanbon(75887)

A bag contains
5 green skittles
3 red skittles
2 orange skittles
6 yellow skittles
how many green skittles must be added to the bag so that the probability of randomly picking a green skittle is 2/3? explain.
---------------------------------------------
Let the number to add be "x":
(5 + x)/(16 + x) = 2/3
Cross multiply to get:

15+3x = 32 + 2x
x = 17
========
Cheers,
Stan H.